Let X be a continuous random variable defined on the interval [0, 4] with probability density function p(x) = c(1 + 4x)

(a) Find the value of c such that p(x) is a valid probability density function.

(b) Find the probability that X is greater than 3.

(c) If X is greater than 1, find the probability X is greater than 2.

(d) What is the probability that X is less than some number a, assuming 0 < a < 4?
Answer #1

a)for this to be valid: \int_{0}^{4} p(x) dx =\int_{0}^{4} c(1+4x) dx= c*(x+2x2) |40 =c*36 =1

c=1/36

b)P(X>3)=\int_{3}^{4} p(x) dx =\int_{3}^{4} (1/36)(1+4x) dx= (1/36)*(x+2x2) |43 =5/12

c)

P(X>2|X>1)= (1/36)*(x+2x2) |42 / (1/36)*(x+2x2) |41 =(13/18)/(11/12)=26/33

d)

P(X<a)=\int_{0}^{a} p(x) dx =\int_{0}^{a} (1/36)(1+4x) dx= (1/36)*(x+2x2) |a0 =(1/36)*(a+2a2)